OpenStudy (anonymous):

inverse of functions f(x)=4(x+1)

OpenStudy (anonymous):

x=4(y+1) x/4=y+1 y=x/4-1 f^-1(x)=x/4-1

OpenStudy (anonymous):

alright thanks